OpenStudy (anonymous):

The length of a rectangle is three more than two times the width. If the perimeter of the rectangle is 134 inches. What is the area of the rectangle.

OpenStudy (anonymous):

A = (w)(2w + 3) 134 = 2w + 2L = 2w + 2[(2w + 3)] From the second equation you can get what "w" is. Then, you can plug that back into equation #1.

OpenStudy (anonymous):

Solve the following equations for L and w.\[\{L=2w+3,2(L+w)=134\} \]\[\left\{L=\frac{137}{3},w=\frac{64}{3}\right\} \]\[\text{Area} = L*w=\frac{137}{3}*\frac{64}{3}=\frac{8768}{9}=974.222 \]